### v1

# Multi-Modal Multi-Task (M3T) Federated Foundation Models for Embodied AI: Potentials and Challenges for Edge Integration

Abstract
As embodied AI systems become increasingly multi-modal, personalized, and interactive, they must learn effectively from diverse sensory inputs, adapt continually to user preferences, and operate safely under resource and privacy constraints. These challenges expose a pressing need for machine learning models capable of swift, context-aware adaptation while balancing model generalization and personalization. Here, two methods emerge as suitable candidates, each offering parts of these capabilities: Foundation Models (FMs) provide a pathway toward generalization across tasks and modalities, whereas Federated Learning (FL) offers the infrastructure for distributed, privacy-preserving model updates and user-level model personalization. However, when used in isolation, each of these approaches falls short of meeting the complex and diverse capability requirements of real-world embodied environments. In this vision paper, we introduce Federated Foundation Models (FFMs) for embodied AI, a new paradigm that unifies the strengths of multi-modal multi-task (M3T) FMs with the privacy-preserving distributed nature of FL, enabling intelligent systems at the wireless edge. We collect critical deployment dimensions of FFMs in embodied AI ecosystems under a unified framework, which we name “EMBODY”: Embodiment heterogeneity, Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, and Yielding safety, privacy, and personalization. For each, we identify concrete challenges and envision actionable research directions. We also present an evaluation framework for deploying FFMs in embodied AI systems, along with the associated trade-offs.
I Introduction
Embodied AI refers to artificial intelligence systems that are physically situated in the world — typically within robots or agents that can sense, act, and learn through interaction with their environment. Building on this foundation, embodied AI is not just redefining the role of intelligent systems, it is reimagining their very nature. What fundamentally distinguishes the next generation of embodied AI agents from traditional AI systems, such as Large Language Models (LLMs) or static vision classifiers, is their demand for interactive, physically-grounded intelligence [1]. In particular, embodied agents must continuously perceive the world through multiple sensor modalities (e.g., vision, touch, audio), interact with dynamic environments, and adapt to diverse tasks, from object manipulation and social interaction to navigating unstructured terrain for search-and-rescue and assisting surgeons in hospitals.
These requirements cannot be met by the deployment of narrowly trained, single-task models at these AI agents. Instead, they naturally align with the capabilities of emerging multi-modal multi-task (M3T) foundation models (FMs), which are large-scale architectures often pretrained on diverse datasets that span language, visual scenes, and human instructions [2]. M3T FMs can provide a unified semantic backbone for embodied agents, enabling them to interpret instructions, understand environments, and plan actions. For example, a kitchen robot could leverage a single M3T FM to recognize ingredients, follow verbal commands, and manipulate utensils, even in settings it has never seen (e.g., through few/zero-shot learning).
Nevertheless, applying FMs to embodied AI introduces new challenges that call for a migration from their conventional centralized training/fine-tuning. In essence, each robot/agent experiences the world through its own embodiment: different sensors, actuators, tasks, and user interactions. Further, these robots/agents operate in decentralized physical environments (e.g., homes, hospitals, and factories), where through their embodiment they accumulate rich, contextual, and privacy-sensitive data (e.g., confidential industrial processes) that cannot be easily pooled/centralized at scale. Subsequently, to truly realize the potential of M3T FMs in these settings, we should move toward cross-embodiment learning, where embodied agents that are inherently data collectors can share, refine, and adapt M3T FMs through decentralized collaborations despite their heterogeneous configurations. Here, Federated Learning (FL) offers a compelling mechanism for this collaboration, enabling distributed agents to share model updates without transmitting raw data, thereby preserving privacy [3, 4].
In this work, we introduce Federated Foundation Models (FFMs) for embodied AI, a natural yet underexplored solution to the challenges and motivations outlined above. Integration of FFMs in this domain creates a new paradigm that brings together the expressive generalization power of FMs with the privacy-preserving, decentralized adaptation/learning capabilities of FL. To give our discussions a unified theme, we identify the most relevant aspects of embodied AI that affect the implementation of FFMs over the network edge under EMBODY dimensions: Embodiment heterogeneity (hardware, sensors, actuators), Modality richness and imbalance, Bandwidth and compute constraints, On-device continual learning, Distributed control and autonomy, Yielding safety, privacy, and personalization. This work is created with the purpose of being a vision paper that both illuminates the transformative potential of integrating FFMs in embodied AI and expose the key challenges arising from such integration. Our contributions are summarized below.
  • We propose an architecture for FFMs suitable for embodied AI, featuring modular sensor encoders, Mixture-of-Experts (MoE) layers, and task heads.
  • We concretize the EMBODY dimensions, showcase the unique capabilities of FFMs for embodied AI, and highlight various use cases of FFMs in this domain.
  • We outline various research directions grounded in the EMBODY dimensions, unveiling the unique opportunities that the modularity of FFMs offer for embodied AI. These directions are intentionally framed to capture the broader theme of “what can be done” in this underexplored research area, offering a flexible conceptual basis for future decomposition into specific, actionable research studies.
  • We envision an evaluation framework for FFMs in embodied AI, consisting of different evaluation metrics and tradeoffs.
TABLE I: Comparative analysis of FL-only, FM-only, and FFM approaches in the embodied AI domain.
Dimension FL-only Robotics Systems FM-only Robotics Systems Federated Foundation Models (FFMs)
Modality Handling Typically uni- or bi-modal (e.g., vision + depth); often fixed during the training Trained on massive cross-modal corpora; supports rich modality fusion Supports rich, time-varying multi-modal input (vision, haptics, and audio) via modular encoders and shared latent representation
Training Setup Typically distributed training from scratch Centralized pretraining; often uses large-scale datasets Decentralized cooperative modular training; enables continual modular modifications
Personalization Local model adaptation or clustering-based personalization Typically lacks per-agent personalization; assumes centralized adaptation Lightweight personalization via on-device modules, prompt tuning, or adapters while preserving global backbone consistency
Generalization Across Embodiments Weak; performance drops across agent types or sensor configurations Moderate; limited adaptability to embodiment diversity and emerging modalities/tasks without centralized retraining/fine-tuning Strong; supports embodiment-conditioned learning, modular control heads, and cross-agent module transfers
Scalability Across Agents Designed for multi-agent participation Not designed for multiple-agent adaptation or training Designed for multi-agent participation with sparse, asynchronous updates and scalable modular coordination mechanisms
Privacy Guarantees Strong; data remains on-device; suitable for sensitive environments Weak; requires centralized data ingestion and fine-tuning Strong; Maintains privacy through federated aggregation while keeping the raw data local
Adaptation Frequency Episodic or periodic retraining Offline adaptation; large update intervals; not suitable for in-the-loop updates Real-time or task-triggered on-device adaptation via modular updates (e.g., encoders, adapters, prompt tuning)
Update Efficiency Lightweight updates due to the use of small models but often low expressiveness Efficient modular, adapter, or prefix/prompt-based updates, centralized fine-tuning required; privacy-risk Efficient modular, adapter, or prefix/prompt-based updates suitable for distributed privacy-preserving edge deployment
Safety and Interpretability Easier to interpret due to model simplicity; can be fine-tuned to obey the safety measures; lacks generalization Hard to interpret; Modular components support targeted auditing and safe rollback of local adaptations; emergent behavior may violate safety unless centrally verified Hard to interpret; Modular components support distributed on-device targeted auditing and safe rollback of local adaptations
Prompt Tuning Support Rarely supported or relevant due to small model scale Centralized prompt tuning enables efficient adaptation to new tasks and sensor conditions without modifying core weights Distributed on-device prompt tuning enables efficient adaptation to new tasks and sensor conditions without modifying core weights
Figure 1: (a): A schematic of FL architecture, where clients/agents engage in collaborative model training. The collaborative training occurs through the repetition of (i) local model training at the agents using their local data, (ii) model/gradient transmission to the server, and (iii) updating the global model at the server based on the received models/gradients (e.g., via weighted averaging) and the broadcast of the global model from the server to the agents to initiate the next round of local model training. (b): Modular architecture of M3T FMs, comprising modality encoders, MoMEs, MoTEs, and task heads. Based on the input characteristics, a subset of MoMEs are triggered/engaged in the inference and training. Further, based on the desired output tasks, a subset of MoTEs will be activated. (c): Architecture of FFMs in embodied AI, where different agents have different modalities and tasks of interest. Each agent possesses a local FM, trains different modules (e.g., encoder, task head, subset of MoMEs or MoTEs) of its local FM, and transmits them to the server for aggregation. The server aggregates the received modules and broadcasts these aggregated modules back to the agents. The received aggregated modules can further go through a local fine-tuning at the agents.
II Background and Related Work
II-A Embodied AI
Modern instances of embodied AI agents are humanoid robots (e.g., Figure 01 by Figure AI), mobile robotic platforms (e.g., Boston Dynamics’ Spot), and extended reality (XR) agents operating through immersive wearables, such as Meta Quest or Apple Vision Pro [1, 5]. Unlike traditional AI systems, mostly confined to screen-based tasks, embodied agents interact with dynamic, uncertain physical and social environments.
Embodied agents fundamentally depend on multi-modal perception to gain an understanding of their surroundings through a diverse range of sensory data, such as visual (RGB, depth), auditory, tactile, inertial, and haptic, often in real-time and under varying quality conditions. Also, to enable goal-directed behaviors, these agents may need to tightly couple perception with motor actions, comprising walking, turning, or navigating (i.e., locomotion), grasping or moving objects (i.e., manipulation), gesturing or signaling (i.e., expressive movement), and operating tools or interfaces. Such motor actions should often be performed under strict latency constraints to respond to environmental or user-driven changes. Nevertheless, unlocking the use of the multi-modal perceived data of these agents to train machine learning (ML) models capable of guiding motor actions under strict latency constraints is not a trivial task: these agents, such as home robots or wearable XR systems, collect deeply personal data (e.g., household layouts, gestures, voice), which cannot be transferred to centralized ML model training units (e.g., cloud servers) due to privacy concerns. This motivates the introduction of privacy-preserving distributed ML realized by FL to embodied AI.
II-B FL in Robotics/Embodied AI
Conventional FL (see Fig. 1(a)) operates through a repeated three-step process until model convergence: (i) each client/device trains a local model using its data; (ii) model updates, such as parameters/gradients, are periodically shared with a server/aggregator; and (iii) the aggregator combines these updates (e.g., via weighted averaging) to a global model and broadcasts it to the devices, thereby synchronizing their local models and initiating the next round of training. FL has been applied in robotics and embodied AI for tasks such as cooperative driving and collaborative manufacturing, enabling robots to jointly learn motion plans and safety-critical controls. Its variants, such as decentralized/gossip-based FL, federated reinforcement learning (FRL), and blockchain-based FL, have also been explored. For example, gossip-based FL has been used in cobot systems, while FRL is used for multi-agent policy learning. For a detailed discussion, see [3, 4]. Collectively, these advances underline FL’s role in scalable intelligence for embodied AI; however, these works do not focus on the FMs.
II-C FMs for Embodied AI
FMs have evolved rapidly, starting with single-modal LLMs (e.g., BERT, GPT-3, and PaLM), showcasing strong text generation/comprehension [6]. This was followed by multi-modal FMs, such as CLIP, DALL-E, and Imagen, enabling zero-shot classification and text-to-image generation [7]. Recently, M3T FMs (e.g., Gato, Kosmos-2, and GPT-4) have emerged, aiming for general-purpose AI that learns/reasons/acts across diverse tasks and modalities using unified architectures [8]. Although FMs are new to embodied AI, some pioneering works exist: RT-1 [9] trained a transformer on 130k robot trajectories; RoboCat [10] explored cross-embodiment learning; and SayCan [11] and ChatGPT-for-Robotics [12] applied LLMs for planning and code generation. These studies show the promise of FMs in embodied AI but assume centralized training and overlook the modularity of modern M3T FMs, an aspect we will explore in the context of FFMs.
II-D FFMs for Distributed Embodied AI
FFMs, especially when considering the emerging M3T FFMs, are a highly recent research topic. Subsequently, they are quite unexplored in embodied AI. Nevertheless, recent research on FFMs in other domains has shown their tremendous potential, and the study of their aggregation methods, computational/communication efficiency, and trustworthiness is gaining substantial interest in the AI/ML and wireless networking community (refer to [13] and the references therein). In this work, we aim to provide one of the first visions for the integration of M3T FFMs in the embodied AI domain. To provide a more structured understanding, in Table I, we provide a comparative analysis of three major paradigms in embodied AI: FL-only systems, FM-only systems, and our proposed FFMs.
II-E Modular Architecture of M3T FFMs
There is no unified architecture for M3T FMs as they are still under active development and envisioned differently across tech companies and academic literature. In this work, we build upon the proposal architecture in [14] and decompose it into a modular architecture, which is depicted in Fig. 1(b), where updates/training can be applied to its various modules independently. Subsequently, to enhance the comprehension of the M3T FFMs, we put them into the context of embodied AI, as illustrated in Fig. 1(c), and explain their components below.
1. Modality Encoders: Each sensory input (e.g., RGB-D images, audio, force/torque signals, inertial readings) is processed through an encoder to transform raw signals into latent representations. This modular integration of encoders enables modality-specific fine-tuning and inter-agent encoder swapping without touching the shared backbone defined below.
2. Shared Backbone: The backbone consists of a set of mixture-of-experts (MoEs) described below.
(i) Mixture-of-Modality Experts (MoMEs): To account for heterogeneity in sensing capabilities and workload balancing, latent representations of modalities are passed through a MoME layer. Experts, which are neural/transformer networks, are selectively activated based on the input characteristics, enabling efficient specialization without full-model activation.
(ii) Mixture-of-Task Experts (MoTEs): To handle the wide range of embodied tasks, such as navigation, object manipulation, gesture following, or environmental interaction, task-specific MoEs are integrated into the model’s pipeline. These allow the model to dynamically activate relevant expert pathways based on a task prompt or contextual signal.
Through the above MoEs, which could be initially pretrained (e.g., through the data scraped from public websites) or trained from scratch alongside other modules [14], the backbone fuses information across modalities and tasks. Also, it captures compositional structure, spatiotemporal relationships, and contextual grounding necessary for embodied AI, while remaining mostly (but not entirely) frozen during deployment.
3. Task Heads: Each embodied task is supported by output heads, which are neural layers that map shared features into concrete predictions (e.g., control commands, action probabilities). This modular integration of heads enables task-specific fine-tuning and inter-agent task head swapping.
4. Adapters and Prompts: Small/shallow adapter modules and/or prompt tuners can be inserted into the backbone or prepended to the input. These modules support agent adaptation or embodiment-specific tuning (i.e., adapting to physical/sensory characteristics of a specific embodied agent).
5. Coordinator and Learning Process: A central coordinator/server manages model updates across agents, following the standard FL process (i.e., local training, aggregation, and broadcast) without exchanging full-model parameters. Here, agents update only local FM sub-components/modules (e.g., encoders, heads, adapters, expert weights, prompts), whose selection can be optimized as we discuss later in the future research directions. The coordinator aggregates these modules and returns them to agents for further updates.
$\star$ Henceforth, we use ‘FFM’ to refer to ‘M3T FFM’ since our focus is solely on this type.
Figure 2: Visualizations of the envisioned future research directions. (a): Embodiment Heterogeneity. Left Plot: Embodiment-Conditioned Expert Routing. Right Plot: Cross-Embodiment Consistency Regularization. (b): Modality Richness and Imbalance. Left Plot: Modality-Aware Clustered Training. Right Plot: Cross-Modal Recovery via Expert Cascades. (c): Bandwidth and Compute Constraints. Left Plot: Context-Aware Federated Caching. Right Plot: Context-Aware and Event-Triggered Scheduling of FFM Modules.
III Use Cases of FFMs in Embodied AI and EMBODY Dimensions
Given the aforementioned notable success of FMs in the embodied AI, FFMs can take this one step further and transform a broad spectrum of embodied AI applications across industrial, domestic, and immersive environments. In the following, we provide some examples and then unveil the natural presence of EMBODY dimensions while articulating them.
Use Cases of FFMs in Embodied AI: We next provide three examples on the use cases of FFMs in embodied AI. (1) FFMs can empower smart factories, where cooperative robots with diverse sensors, actuators, and policies adapt to dynamic workflows and human collaborators. Unlike traditional FMs that may need centralized retraining, FFMs enable decentralized training/adaptation, preserving proprietary data and meeting strict privacy regulations. (2) In domestic environments, FFMs enable assistive robots to learn and personalize to users’ routines, preferences, and spaces without exposing sensitive household data, overcoming a key limitation of FM-only solutions. (3) In immersive XR systems, FFMs enable distributed headsets and wearables to collaboratively fine-tune models for body-language, gaze, and gesture recognition, achieving real-time on-device personalization.
Unlike conventional FMs, which are frozen after centralized pretraining and often lack situational responsiveness, FFMs unlock continuous, privacy-preserving adaptation to new users, tasks, and hardware configurations in all the above scenarios.
Articulation of EMBODY Dimensions: In the above scenarios, the EMBODY dimensions naturally emerge, underscoring the critical challenges that naive implementations of FFMs in embodied AI cannot address.
  • Embodiment heterogeneity spans all the aforementioned scenarios: in smart factories, robots differ in morphology, sensing, and actuation; in homes, assistive robots must adapt to user-specific layouts and hardware configurations; and in XR systems, headsets and embodied avatars vary in tracking precision, interface latency, and sensor fidelity.
  • Modality richness and imbalance is also central across scenarios: factory robots are exposed to vision, force feedback, and machine states; domestic agents interpret multi-modal cues like voice, touch, and gaze; and XR systems may rely on the fusion of head pose, hand motion, eye-tracking, and speech.
  • Bandwidth and compute constraints are common across scenarios: factory robots operate with limited communication windows and rely on on-board processors that cannot support large-scale model retraining; domestic robots are often low-power, cost-sensitive devices lacking powerful GPUs for full-model fine-tuning; XR systems demand ultra-low latency and high frame rates, precluding heavy model updates or large-scale communication during the execution of XR tasks.
  • On-device continual learning manifests itself universally: robotic arms in factories must adapt to new tasks or tools, home assistants to changing user routines, and XR avatars to evolving behavioral signals, all requiring (near) real-time local model/behavior updates without server interventions.
  • Distributed control and autonomy is intrinsic to these scenarios: factory robots may need to coordinate actions without frequent centralized commands, household robots often operate semi-independently across rooms or homes, and XR users can move independently in immersive environments.
  • Yielding safety, privacy, and personalization is a shared imperative across the scenarios: safety/regulatory compliance in factories, user privacy in domestic settings, and individualized experiences in immersive systems all demand personalized intelligence that respects privacy and operates within safety margins.
IV Open Research Directions: An EMBODY-Aligned Agenda for FFMs
We next revisit the aforementioned EMBODY dimensions, aiming to tailor a series of research directions to address them. Each part introduces a core challenge and proposes open research directions grounded in the modular FFM framework. The directions are intentionally framed at a high-level to allow for diverse interpretations and encourage innovation in this underexplored domain.
Figure 3: Visualizations of the envisioned future research directions. (a): On-Device Continual Learning. Left Plot: Drift-Aware Module Aggregation. Right Plot: Temperature-Aware Model Aggregations. (b): Distributed Control and Autonomy. Left Plot: Federated Multi-Agent Contrastive Alignment. Right Plot: Decentralized MoE Federation. (c): Yielding Safety, Privacy, and Personalization. Left Plot: Module-Aware Differential Privacy. Right Plot: Federated Behavioral Shadowing Buffers.
IV-A Embodiment Heterogeneity
The embodied heterogeneity, caused by agents operating in various physical environments, poses a challenge to building shared FFMs that generalize across embodiments (e.g., can be trained on one or more types of robots and still perform effectively when deployed on others). To address this, we envision the following (see Fig. 2(a)):
(i) Embodiment-Conditioned Expert Routing: Modality-specific MoME layers can be augmented with learned embodiment tokens that encode hardware traits (e.g., kinematic range, actuator count, sensor precision). These tokens can then act as soft keys for expert selection, enabling information routing decisions inside the model that reflect physical context during inference and model personalization.
(ii) Cross-Embodiment Consistency Regularization: During federated training, agents with different embodiments often diverge in representation space (i.e., variations in hardware, sensing, or control may lead to different internal encodings of similar tasks). To address this, we envision embedding a consistency loss between related tasks performed by heterogeneous embodiments to make a closer connection between similar tasks performed across embodiment shifts (i.e., changes in the physical configuration of agents, such as differences in arm length, gripper type, or sensor placement, which affect how the same task is perceived and executed).
IV-B Modality Richness and Imbalance
To handle modality richness and imbalance of agents, caused by the (temporal) variations of their sensory inputs, we envision the following (see Fig. 2(b)):
(i) Modality-Aware Clustered Training: We envision agent clustering techniques that group agents with similar tasks while ensuring the combined modalities in each group/cluster cover broader modalities than any single agent. Within each cluster, the server can perform federated aggregation across selected MoME experts, enabling agents to learn not only from peers with similar tasks but also from complementary modalities they lack. Crucially, in each cluster, experts within the MoME layer of one agent can also learn the hidden structures and relationships between modalities by leveraging knowledge from agents with richer modality sets.
(ii) Cross-Modal Recovery via Expert Cascades: To improve robustness against missing modalities, we envision expert cascades within the MoME layer: a series of specialized models that activate only when certain types of data are missing or low-quality. These experts should be trained in a federated manner, where agents locally train their experts using their available sensor data and send only the expert model updates to the server. The server can then aggregate these updates for each of these experts across agents, allowing knowledge sharing without exposing raw data. This way, agents can collaboratively learn to fill in missing modalities (e.g., predicting touch input from vision and movement sensors) while keeping the training decentralized and privacy-preserving.
IV-C Bandwidth and Compute Constraints
To consider the fact that many embodied agents operate on edge processors with strict energy/compute and bandwidth limitations, we envision the following (see Fig. 2(c)) :
(i) Context-Aware Federated Caching: We envision a context-aware edge caching mechanism for FFMs, where the edge server caches the modules received from the agents during the model aggregations, adhering to its memory constraints. Here, agents participating in FFM training first query their needed modules from their nearest server based on their current task and environment (e.g., “following the user in hallway”) to avoid redundant local model updates and transmissions. If another agent has contributed a similar module to the edge cache during previous federated rounds, the querying agent can download and reuse that module with only a light fine-tuning. This mechanism can leverage semantic task signatures (e.g., compact embeddings from recent observation-action histories) to match cached modules with the agent’s current context.
(ii) Event-Triggered Scheduling of FFM Modules: In FFMs, we envision that event-triggered scheduling can reduce unnecessary computation/communication. Specifically, rather than relying on fixed periodic model transmissions to the server, agents can dynamically trigger the training of specific FFM modules (e.g., encoders, adapters, or expert layers) only when significant context/environment changes, task shifts, or performance drops are detected. For example, a mobile robot might update its navigation task head only when entering a new environment or encountering unfamiliar obstacles.
IV-D On-Device Continual Learning
To handle on-device continual learning of embodied agents, caused by facing evolving environments, user preferences, and task definitions — see Fig. 3(a) — we envision the following:
(i) Drift-Aware Module Creation: We envision task drift management for FFMs. At the agent-level, task drift may indicate the need to fine-tune or split a local task head into two. For example, in the case of “cleaning” task for a home robot, given the variations in room layouts, it may benefit from separate heads for “kitchen cleaning” and “bedroom cleaning.” The key challenge lies in determining whether this drift is isolated/local or systematic/global. If drift patterns are detected across multiple agents (e.g., via cluster-level model divergence metrics), the server may create a new global task head. Otherwise, agent-level fine-tuning (e.g., cross-task proximal terms or adding adapters) remains the preferred path to avoid global model bloat and maintain its generalization.
(ii) Temperature-Aware Model Aggregations: In dynamic FL environments, where data and modality drift are common, different model modules can be at varying levels of readiness or maturity — some are well-trained and performing reliably, called “warm”; while some are newly introduced or still adapting (e.g., a freshly added encoder for an emerging modality), called “cold”. To account for these variations, we envision a temperature-aware aggregation strategy, where the FL server adjusts the weight or frequency of modules’ aggregation based on their current “temperature.” Well-performing, stable modules can be sent for aggregation less frequently to avoid marginal performance gains, while newer or less confident modules receive more frequent and aggressive aggregation to accelerate their performance across agents.
IV-E Distributed Control and Autonomy
To enable embodied agents to adapt to their local tasks and contexts in a coordinated yet decentralized manner — see Fig. 3(b) — we envision the following:
(i) Federated Multi-Agent Contrastive Alignment: Embodied agents often can perform similar tasks while operating in different environments, using different hardware, or facing different user preferences. We envision a shared contrastive learning objective that works through teaching agents to pull the internal representations (latent spaces) of similar tasks closer and push apart those of non-similar tasks. In our case, this will align the task understanding of agents with similar tasks without sharing raw data. For example, an aerial drone and a ground robot both performing obstacle avoidance would learn to make their internal representations of this task similar, even though their sensors and actions differ.
(ii) Decentralized MoE Federation: Instead of aggregating entire models, agents can share task-specific MoME/MoTM/encoders/heads via peer-to-peer (P2P) communications (i.e., rather than uploading the full model to a central server, agents directly exchange only the relevant components tied to specific tasks, such as navigation or manipulation, enabling decentralized collaboration and localized knowledge transfer).
IV-F Yielding Safety, Privacy, and Personalization
To make agents comply with safety, privacy, and user-specific constraints — see Fig. 3(c) — we envision the following:
(i) Module-Aware Differential Privacy: To prevent data leakage (e.g., via data inference attacks applied on the transmitted model updates from agents to the server), we envision module-aware differential privacy: a more fine-grained approach to privacy in FFMs, which applies non-uniform noise across different modules, such as encoders, heads, or MoME/MoTEs, before agents transmit them for aggregation. In this approach, the level of noise can be adjusted based on each module’s importance and usage pattern. For example, an encoder capturing generalizable features, or a frequently activated expert for sensitive tasks, would receive stronger noise, while task-specific heads or rarely used experts would require less, as they pose lower privacy risks.
(ii) Federated Behavioral Shadow Buffers: To ensure that model updates from the server are safely tested before being used by agents, we envision first deploying these updates in a “shadow mode” instead of immediately applying them, where the updated model makes predictions but its actions are not executed. The agent compares these predictions to those from its old model and looks for any big differences, while also paying attention to user reactions (e.g., corrections or feedback). This helps the agent and the system check if the new updates improve performance without causing safety issues or unwanted behavior. This shadowing process can be personalized for each user or task and activated only in sensitive situations, such as when robots take actions in shared spaces with people. Note that another option is to test the new model’s actions in a virtual environment, e.g., a digital twin, to safely observe the model’s action consequences without real-world risks.
TABLE II: Key trade-offs in FMM design for embodied AI (framed under EMBODY Dimensions). Letters ‘E’, ‘M’, ‘B’, ‘O’, ‘D’, and ‘Y’ in the last column refer to various dimensions of EMBODY.
Trade-off Dimension Option A (Advantage) Option B (Advantage) Key Tension / Real-World Implication EMBODY Dimension
Personalization vs. Global Consistency Strong per-agent adaptation via fine-tuning on-device modules (adapters, prompts, experts) Stable cross-agent coordination through shared frozen backbones Over-personalization may cause over-fitting; overly generic models may underfit unique agent dynamics E, M, O, D, Y
Communication Frequency vs. Model Quality Infrequent updates save energy and bandwidth Frequent updates improve convergence and model quality Infrequent updates can lead to low model quality; frequent updates lead to a higher resource overhead M, B, O
Privacy Assurance vs. Model Quality Strong privacy via adding a high differential privacy (DP) noise Full gradient/data sharing without adding DP noise, leading to a higher quality model Privacy constraints may suppress the model quality; Full gradient sharing risks data inference attacks M, Y
Low-overhead Modular Updates vs. Extensive Fine-tuning Light updates of the adapters, subset of MoEs, and prompts enable low-cost local tuning Extensive updates of the adapters, subset of MoEs, and prompts unlock a better learning capacity Light modular tuning aids edge deployment over resource-constrained agents but may limit the model performance E, M, B, O, Y
Safety Guarantees vs. Exploration Agility Risk-averse model updates reduce unsafe actions and enable certification Fast adaptation without risk-considerations improves exploration and model agility Safety conservatism may stall exploration; aggressive updates may introduce unsafe behavior O, D, Y
TABLE III: Benchmarking FFMs with existing embodied AI datasets: Dataset characteristics and their offered modalities and tasks [15].
Dataset / Simulator Supported Tasks Benchmark Modality Types FFM Evaluation Goals
Habitat Navigation, object search, instruction following RGB, depth, egomotion Evaluate cross-embodiment generalization, modular task head updates, zero-shot inference
Meta-World Robotic manipulation RGB, proprioception, joint Force/Torque Assess leave-one-embodiment-out generalization, task head reuse, modular fine-tuning
ManiSkill Dextrous Hand, Mobile Manipulation RGB, depth, proprioception, Egomotion Test robustness to missing/noisy modalities, cross-modal recovery via expert cascades
iGibson / RoboSuite Object Manipulation, Navigation, Household Activities RGB, depth, semantic segmentation Context-aware module scheduling, evaluation of safety-aware updates and behavioral shadowing
BEHAVIOR-1K cleaning tasks, cooking activities, organizational tasks Visual inputs, semantic segmentation, object states Benchmark continual learning and adaptation across diverse goals and environments
V Evaluation Dimensions of FFMs: An EMBODY-Aligned Perspective
We next envision an evaluation framework developed around the EMBODY dimensions.
1. Embodiment Heterogeneity: To assess generalization across diverse robot morphologies and sensor configurations, models should be evaluated under “leave-one-embodiment-out” settings. This involves training on $N-1$ embodiments and deploying zero-shot on the $N^{\text{th}}$. Key metrics encompass task success rates, policy stability under various morphologies, and safety violations. Benchmark environments should simulate variable kinematics, sensor fidelity, and actuation delay.
2. Modality Richness and Imbalance: FFMs should be robust to missing or sparse modalities across agents. Evaluations should thus measure performance degradation under modality dropout, imbalance, or late-fusion timing mismatches (i.e., cases where different sensory modalities, such as vision, touch, or proprioception, arrive asynchronously or fused with misaligned timestamps due to differing sensor update rates or processing delays). To evaluate model’s robustness, the following metrics can be envisioned: modality ablation accuracy (i.e., the model’s performance when one or more input modalities are intentionally removed/masked), cross-modal transfer ability (i.e., the extent to which knowledge learned from one modality, such as vision, can be used to infer or compensate for another, such as audio or touch), and reliance bias (i.e., over-dependence on a single modality).
3. Bandwidth and Compute Constraints: FFMs deployed on resource-constrained agents must be assessed for communication and energy efficiency, under various choices of modular training/aggregation (e.g., varying numbers of involved experts, transmitting encoders vs task heads vs experts for aggregation). Evaluation should log model update size, compute per-task completion, energy-to-performance ratio, model aggregation latency, and processing depth-related metrics (e.g., percentage of activated experts during model updates and execution).
4. On-Device Continual Learning: Continual learning capability of the FFMs should be evaluated by measuring the model adaptation speed and rates of knowledge forgetting over changing environments. Metrics include forward and backward transfer (i.e., the ability to leverage prior knowledge to learn new tasks, and the ability to retain old knowledge after learning new tasks), time-to-adapt (i.e., how quickly an FFM improves performance on a new task/environment), and retention curves (i.e., performance trajectories that track how well previously learned tasks are maintained over time in the model).
5. Distributed Control and Autonomy: FFMs should be evaluated for coordination effectiveness, latency of global/local model updates, and cross-agent generalization. Metrics may include behavior divergence across agents (i.e., the degree to which different agents trained under the same FFM develop inconsistent or conflicting behaviors), emergent behavior alignment (i.e., how well the trained agents converge on coordinated or complementary behaviors during interaction with their environments), and robustness to asynchronous updates (i.e., the model’s ability to maintain stable performance when agents receive and apply model updates at different times).
6. Yielding Safety, Privacy, and Personalization: Evaluation must explicitly capture FFMs’ safety violations, such as collisions, misclassifications of dangerous zones, or regulatory compliance lapses. Also, privacy-preserving utility should be measured by how techniques such as differential privacy applied across various model modules can affect downstream embodied AI task performance. Further, personalization should be assessed in terms of task performance gain per user/agent (i.e., how much an individual agent’s task performance improves after local adaptation) and misalignment risks (i.e., the potential for personalized updates to diverge from global objectives or introduce unintended behaviors).
Design Tradeoffs and Benchmarking: To operationalize FFMs in embodied AI, it is critical to understand the multi-dimensional trade-offs that emerge when balancing model personalization capability, cross-agent consistency, resource constraints, safety, and interoperability. To facilitate comprehension, in Table II we present a structured view of these trade-offs, framed through the lens of the EMBODY dimensions. Each row highlights a tension between two competing design choices, such as personalization vs. global generalization, or modular updates vs. full fine-tuning, and articulates their real-world implications. This table can serve as a guide for system designers to tailor FFM architectures and protocols according to the operational priorities and constraints of their target environments. Further, in Table III we provide a structured description of key embodied AI benchmarks that can be used to evaluate FFMs across diverse tasks and modalities.
VI Conclusion
In this paper, we envisioned that by integrating the generalization power of M3T FMs with the decentralized learning capabilities of FL, FFMs can offer a unified learning approach in the embodied AI ecosystems. Through the EMBODY framework, we articulated the core dimensions that should be addressed for FFMs to succeed in real-world embodied AI deployments. For each dimension, we identified open challenges and envisioned actionable research directions that span model architecture, training dynamics, and system configuration. We also discussed evaluation protocols and trade-off analysis for FFM deployment over embodied AI.